OpenStudy (anonymous):

A car 4m long came up behind a semi-trailer 20m long travelling at a steady 72km/h. After travelling behind the semi-trailer for some time, the driver of the car decided to overtake. If the car pulled out from behind the semi-trailer with the front of the car 10m behind the rear of the trailer and accelerated at 3.5m/s/s until the rear of the car was 10m in front of the truck, how far would the car travel and how long would it take? I used 44m as displacement, and got velocity from that, which I then used to get time. But my answers are wrong. Any ideas?

OpenStudy (anonymous):

how did you get 44m as a displacement?

OpenStudy (anonymous):

The whole overtaking process. 4 + 10 + 20 + 10

OpenStudy (anonymous):

ok then I think you are looking for s=v0t+(at^2)/2 where v0=initial velocity a=acceleration t=time elapsed and s=distance traveled

OpenStudy (anonymous):

I don't have any value for t though.

OpenStudy (anonymous):

you are solving for t, the initial velocity is zero as well. The really questions is how long does it take to travel 44m is accelerating at 3.5 m/s Both the car and truck are moving at the same speed, at the start, the 72km/h doesn't matter

OpenStudy (anonymous):

Solving that equation for t has never really worked for me. Would t = v-u/a work?

OpenStudy (anonymous):

44=(3.5t^2)/2 88=(3.5t^2) 88/3.5 is about 25, so the answer is about 5 sec

OpenStudy (anonymous):

Acceleration is a "compounding" effect, so the speed, curves and there the power of 2 if that make any sense to you

OpenStudy (anonymous):

You're right. The correct answer is 5. But I'm still a little confused, what formula was that using? And that kind of makes sense, but I'm not sure how that fits into the equation..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

try the formula S=ut+1/2a(t)square

OpenStudy (anonymous):

Been there, done that.

OpenStudy (anonymous):

 suppose car overtakes in time t, then total distance covered by DRIVER OF TRAILER=20t +we add the legth of trailer as only driver and front are moving next cars driver moves 10m+10m+20t+20m+4m(its own length) then u would have s=1/2at^2 formula to use the thiung is that consider the drivers of car and truck next consider the rest of the portion of vehicles